OpenStudy (anonymous):

Which of the following is a solution to 5 square root of x minus 7 equals 10

jimthompson5910 (jim_thompson5910):

\[\Large 5\sqrt{x} - 7 = 10\] \[\Large 5\sqrt{x} = 10+7\] \[\Large 5\sqrt{x} = 17\] \[\Large \sqrt{x} = \frac{17}{5}\] \[\Large x = \left(\frac{17}{5}\right)^2\] \[\Large x = \frac{289}{25}\]